OpenStudy (anonymous):

what affect does h = -2 have on the parent function y=1/x ?

OpenStudy (anonymous):

Graph moves 2 units up Vertical stretch by a factor of |-2| Graph moves 2 units down Graph moves 2 units to the left

OpenStudy (cwrw238):

y = (1/x ) - 2 moves graph 2 units down

OpenStudy (anonymous):

thanks! do you know: What affect does k = 5 have on the parent function y=1/x ?

OpenStudy (anonymous):

Graph moves 5 units to the right Vertical stretch by a factor of |5| Graph moves 5 units up Graph moves 5 units left

OpenStudy (cwrw238):

5 is +5 so 5 units up

OpenStudy (anonymous):

your awesome :)

OpenStudy (anonymous):

what are h and k u talking about..for graph its generally y=f(x) if u increase x , graph shrinks, decrease x graph expands, add something like y=f(x) + c graph shifts on y aaxis by c units